OpenStudy (anonymous):

What is the difference of 3,789 and 705?

OpenStudy (anonymous):

Just subtract.

OpenStudy (anonymous):

i thought so

OpenStudy (anonymous):

9-5=4 8-0=8 7-7=0 3-0=3